OpenStudy (anonymous):

f(x)=6x+3 g(x)=2x^2+1 f(g(x)) is what, and what is the domain? please help :)

OpenStudy (anonymous):

these are polynomials so domain is all real numbers

OpenStudy (anonymous):

\[f(g(x))=f(2x^2+1)=6(2x^2+1)+3\]

OpenStudy (anonymous):

you can of course clean the last one up with some algebra

OpenStudy (anonymous):

ok so if one of the questions is (f/g)(x), what would I do?